The gauge pressure in your car tires is 2.80 ✕ 105 N/m2 at a temperature of 35.0°C when you drive it onto a ship in Los Angeles to be sent to Alaska. What is their gauge pressure (in atm) later, when their temperature has dropped to

−39.0°C?

Assume the tires have not gained or lost any air.
Answer #1

By Ideal gos low PV = nRT PV = nR constant Pivi Ti PEVE Te Vi = 14 Pi - PE Meht Multiple each side by TePi = 2.8810 Nim? Ti = 35+273=308k = -39+273 =234k Of = 2:8 X10 X234 308 308 X 234 1 0 = 2:12X60% atm]